OpenStudy (skypie13):

Chris has a resting heart rate of 68 beats per minute. He is 13 years old. Calculate Chris’ Target Heart Rate Zone and show your calculations. What should Chris do if his heart rate becomes higher or lower than the limits of his Target Heart Rate Zone while he exercises?

OpenStudy (mathstudent55):

Do you know how to calculate the target heart rate?

OpenStudy (skypie13):

No

OpenStudy (mathstudent55):

First you need the maximum heart rate (MHR). MHR = 220 - age What is Chris' MHR?

OpenStudy (skypie13):

?????????

OpenStudy (mathstudent55):

What don't you understand?

OpenStudy (skypie13):

Everything. JK. I just don't understand this at all. All I need is the answer....

OpenStudy (mathstudent55):

The maximum heart rate is abbreviated MHR. To find the MHR, just subtract 220 minus the age of the person. Chris is 13 years old. Chris' MHR is 220 - his age. What is 220 - 13?

OpenStudy (skypie13):

207.

OpenStudy (mathstudent55):

Good. Now we do the next step.

OpenStudy (skypie13):

Wazzat?

OpenStudy (mathstudent55):

Now we know two things. Chris' resting hear rate (RHR) is 68 since the problem tells us that. Chris' maximum hear rate (MHR) is 207 since we calculated it.

OpenStudy (mathstudent55):

Now we need to calculate his Target Heart Rate Zone. The target heart rate zone is a range of numbers from a lower heart beat number to a higher heart beat number. We need to calculate both of those numbers. We need to calculate the lower limit and the upper limit of his target heart rate zone. There is a formula for each one.

OpenStudy (skypie13):

You mean addition or subtraction?

OpenStudy (mathstudent55):

Addition, subtraction and multiplication. Here is the lower limit formula. Target Heart Rate zone lower limit: lower limit = MHR – RHR * 0.50 + RHR First, we replace MHR and RHR with our actual numbers. Remember we have MHR = 207, and RHR = 68 lower limit = MHR – RHR * 0.50 + RHR lower limit = 207 – 68 * 0.50 + 68 Now we need to do the calculation above. You must follow the correct order of operations. Do the multiplication first. What is 0.5 * 68?

OpenStudy (skypie13):

68.5. I am a sixth grader. I know how to do math.

OpenStudy (mathstudent55):

Before we go on, do you have a book for this subject? Does your book have its own formulas for these calculations? I noticed that there are different formulas for these numbers. We should use your formulas, so your answer is what your teacher expects.

OpenStudy (skypie13):

I do FLVS. We have no books, only computers. That is why I rely so much on this site.

OpenStudy (mathstudent55):

OK. We'll use the formulas I have then. I'm sure you know math, but you need to be careful. 0.5 * 68 means 0.5 times 68 and it equals 34, not 68.5. You added the numbers. We need to multiply them.

OpenStudy (skypie13):

Sorry, messed up.

OpenStudy (mathstudent55):

No problem. Now that we know 0.5 * 68 = 34, we replace 0.5 * 68 with 64 in the formula. lower limit = 207 – 34 + 68

OpenStudy (mathstudent55):

Now do 207 minus 34 then plus 68. What do you get?

OpenStudy (skypie13):

241

OpenStudy (mathstudent55):

207 - 34 + 68 = 241 Now we use the formula for the upper limit upper limit = MHR – RHR * 0.85 + RHR upper limit = 207 - 68 * 0.85 + 68 upper limit = 207 - 217.2 We get a lower limit of 241 and and upper limit of 217 The lower limit must be a lower number than the upper limit. That means our to formulas were switched. Chris' target heart rate zone is from 217 to 241. .